# AETAS: Analysis of Evolving Temporal Affect and Semantics for Legal History

### v1

@@ 4.7 Split-half stability
Splitting each bin (1750s, 1850s, 1900s) into halves, training SGNS on each half (20 repeats), and measuring within-bin drift shows substantial sampling variance: justice mean split-half drift 0.71 (std up to 0.10), crime 0.63 (std 0.08), transportation 0.72 (std 0.08) (Figure 6(a)). We therefore report a net drift (Figure 7) that subtracts the average split-half baseline and a standardized effect size (net divided by pooled split-half std); only charity (1750s–1900s) remains positive, with a small .
@@ 4.8 Frequency controls and preprocessing
Regressing SGNS drift on mean and delta frequency across spans yields no significant coefficients (p0.09, ); contextual APD and incremental SGNS show similarly non-significant frequency effects (p0.14, ), suggesting observed drift is not explained by frequency trends alone. A lightweight NER/POS filter on sampled texts removes 10–16% of tokens and 21–35% of types (Figure 6(b)), indicating named entities and numerals contribute noticeable noise. Preprocessing: lowercase, strip punctuation except hyphens, no stopword removal, no lemmatization in the main runs.
@@ 4.9 NER/POS filtering sanity check
Applying a lightweight NER/POS filter (spaCy en_core_web_sm) to 200-sample texts per bin removes 10–16% of tokens and 21–35% of types. This suggests that proper-noun/number noise is non-trivial; integrating NER/POS-aware preprocessing is left for future modeling runs.
@@ 5 Discussion
The results align qualitatively with digital-humanities narratives: (1) erosion of transportation in British penal practice; (2) professionalisation of judges and a move from status-based justice to codified procedure; (3) medicalisation of madness and insanity. Yet split-half instability is large, and net drifts after subtracting that baseline are mostly small or negative (Figure 7), so these findings should be read as hypothesis-generating rather than definitive evidence of semantic change.
Robustness additions partially address instability: expanding reduces but does not eliminate neighborhood sparsity (Figure 4); seed-variation bands suggest the mercy-retribution sign of justice is robust (Figure 5); corpus and frequency diagnostics guard against coverage/frequency artifacts; net-drift/z heatmaps contextualize effect sizes (Figure 7). Remaining risks include anchoring bias from using 1900s as the sole base, limited contextual baselines in sparse bins, uncalibrated value axes (no external labels), and proper-noun noise.
@@ 6 Limitations and future work
Bootstrapped stability for drift and axes; hubness diagnostics; token-based contextual baselines with balanced sampling/PRT.
Calibrated value axes with external labels (e.g., sentencing outcomes or expert ratings), additional moral/procedural axes, richer seed bootstraps, and integrating NER/POS filtering into main runs.
Stronger polysemy/context-diversity controls following [7].
Evaluating semantic change remains challenging without gold standards [3]; crowdsourced historical judgments or links to sentencing outcomes would strengthen validation and the “expert system” claims.
@@ 7 Conclusion
We demonstrated an auditable semantic-drift pipeline that couples diachronic embeddings with digital-humanities storytelling. Each module is an inference component whose outputs can be inspected and recombined. Given the high within-bin variance and modest net effects, the current claims should be viewed as hypotheses; with stronger calibration (chain alignments, contextual baselines, calibrated value axes, and human/external validation), the approach can mature into a reusable template for domain-specific semantic change analysis.

### v2

@@ 4.8 Split-half stability
Splitting each bin (1750s, 1850s, 1900s) into halves, training SGNS on each half (20 repeats), and aligning the half models before comparison yields modest within-bin drift (bin-level means 0.06, 0.09, 0.08 respectively; Fig. 7). Net drift after subtracting these baselines remains positive across target terms, with the largest net shifts in justice and transportation (Fig. 9). Standardized effect sizes are large because split-half variance is small; we therefore interpret as a stability diagnostic rather than a strict statistical test. Table 3 provides the full net-drift and values. Using full-corpus split-half variance, per-target 95% intervals for net drift remain positive for all focal words (e.g., justice 0.478–0.521 for 1750s–1850s), and are reported alongside released artifacts.
@@ 4.9 Frequency controls and preprocessing
Regressing SGNS drift on mean and delta frequency across spans yields no significant coefficients (p0.09) on a small sample (27 word-span pairs; ), so we treat frequency controls as descriptive rather than confirmatory; contextual APD and incremental SGNS show similarly non-significant frequency effects (p0.14, ). Partial correlations are weak: drift vs. log mean controlling log delta yields , while drift vs. log delta controlling log mean yields (). A lightweight NER/POS filter on sampled texts removes 10–16% of tokens and 21–35% of types (Fig. 8), indicating named entities and numerals contribute noticeable noise. We additionally run a full case-aware proper-noun proxy (capitalized non-initial tokens, numerals, honorific patterns) on the main bins and report its drift/overlap stability in the robustness section. A conservative orthography normalization (plural-only lemmatisation plus British/American variant mapping) reduces vocabulary by 6–7% on 2,000-document samples; drift rankings correlate strongly with the unnormalized sample (Spearman , ), and justice/transportation remain the highest-drift terms among those retained. We therefore treat normalization as a qualitative check rather than a replacement for the main pipeline. Preprocessing for the baseline runs: lowercase, strip punctuation except hyphens, no stopword removal, no lemmatization.
@@ 4.10 NER/POS filtering sanity check
Applying a lightweight NER/POS filter (spaCy en_core_web_sm) to 200-sample texts per bin removes 10–16% of tokens and 21–35% of types, indicating non-trivial proper-noun/number noise. To scale to the full corpus, we run a case-aware proxy filter (dropping capitalised non-initial tokens, numerals, and honorific patterns) on the main bins and rerun the pipeline. Drift rankings remain highly correlated with the baseline (Spearman , ), with slight overlap gains (mean Jaccard 0.037 for , 0.040 for ), suggesting that named-entity noise is not the primary driver. Sampled POS tagging of occurrences shows justice shifting toward PROPN usage (PROPN share 0.56 in 1750s, 1.00 in 1850s, 0.91 in 1900s), consistent with the title-driven shift.
| Word | Span | Drift | Split | Net | |
| justice | 1750s-1850s | 0.565 | 0.066 | 0.499 | 45.7 |
| justice | 1750s-1900s | 0.520 | 0.069 | 0.450 | 46.8 |
| justice | 1850s-1900s | 0.311 | 0.071 | 0.240 | 20.4 |
| crime | 1750s-1850s | 0.344 | 0.079 | 0.265 | 15.1 |
| crime | 1750s-1900s | 0.388 | 0.077 | 0.311 | 20.7 |
| crime | 1850s-1900s | 0.275 | 0.081 | 0.193 | 11.2 |
| poverty | 1750s-1850s | 0.342 | 0.082 | 0.260 | 12.4 |
| poverty | 1750s-1900s | 0.259 | 0.062 | 0.197 | 12.6 |
| poverty | 1850s-1900s | 0.294 | 0.088 | 0.206 | 10.8 |
| insanity | 1750s-1850s | 0.460 | 0.065 | 0.395 | 27.2 |
| insanity | 1750s-1900s | 0.427 | 0.055 | 0.373 | 25.6 |
| insanity | 1850s-1900s | 0.284 | 0.077 | 0.207 | 14.2 |
| charity | 1750s-1850s | 0.385 | 0.100 | 0.284 | 12.1 |
| charity | 1750s-1900s | 0.500 | 0.098 | 0.402 | 17.3 |
| charity | 1850s-1900s | 0.401 | 0.096 | 0.305 | 13.3 |
| transportation | 1750s-1850s | 0.533 | 0.060 | 0.473 | 42.5 |
| transportation | 1750s-1900s | – | – | – | – |
| transportation | 1850s-1900s | – | – | – | – |
| punishment | 1750s-1850s | 0.373 | 0.070 | 0.303 | 23.7 |
| punishment | 1750s-1900s | 0.349 | 0.060 | 0.289 | 23.8 |
| punishment | 1850s-1900s | 0.313 | 0.086 | 0.228 | 13.2 |
@@ 5 Discussion
The results align qualitatively with digital-humanities narratives: (1) erosion of transportation in British penal practice; (2) professionalisation of judges and a move from status-based justice to codified procedure; (3) medicalisation of madness and insanity. These shifts also coincide with known reform periods (transportation expansion and debate in the eighteenth century, penal-servitude reforms in the mid-nineteenth century, and Victorian-era medicalisation), providing lightweight event-based corroboration. A corpus-wide title filter indicates that professional honorific usage contributes to the justice shift but does not remove it, and Global Anchors drift agrees with the SGNS ranking. After alignment-aware split-half baselines, within-bin variance is modest and net drifts remain positive across terms (Fig. 9). The findings are therefore more stable than the uncorrected split-half results, but they remain hypothesis-generating without stronger external validation.
Robustness additions partially address instability: expanding reduces but does not eliminate neighborhood sparsity (Fig. 4); seed-variation bands suggest the mercy-retribution sign of justice is robust (Fig. 6); corpus, anchor, and hyperparameter diagnostics guard against coverage/frequency artifacts; net-drift/z heatmaps contextualize effect sizes (Fig. 9). Remaining risks include some anchoring bias despite stable-anchor checks, limited contextual baselines in sparse bins, and uncalibrated value axes (no external labels).
@@ 6 Limitations and future work
Richer alignment-free usage measures (e.g., AMD/SAMD-style usage correspondence) and dependency-slot baselines to complement Global Anchors and embedding drift.
@@
Historically adapted contextual models (continued pretraining on OBC, HistBERT-style) and larger context samples for APD/JSD baselines.
Document-level bootstrap stability for drift and axes beyond anchor sampling; hubness diagnostics; token-based contextual baselines with balanced sampling/PRT.
Calibrated value axes with external labels (e.g., sentencing outcomes or expert ratings), additional moral/procedural axes, and richer seed bootstraps.
Stronger polysemy/context-diversity controls following (kutuzov2022).
Evaluating semantic change remains challenging without gold standards (tahmasebi2021); crowdsourced historical judgments or links to sentencing outcomes would strengthen validation and the “expert system” claims.
@@ 7 Conclusion
We demonstrated an auditable semantic-drift pipeline that couples diachronic embeddings with digital-humanities storytelling. Each module is an inference component whose outputs can be inspected and recombined. With alignment-aware split-half baselines, stable-anchor checks, hyperparameter perturbations, and preprocessing ablations, within-bin variance is modest and the main shifts remain robust across focal terms; external validation (contextual baselines, calibrated value axes, event-aligned benchmarks) would further strengthen claims of semantic change.
Data availability
The Old Bailey Corpus is distributed under its own license and cannot be redistributed here. We provide derived drift tables, figures, and scripts to reconstruct results for licensed users.
Funding
This work received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
AI Disclosure Statement
The author used OpenAI Codex to assist with code drafting and refactoring for the analysis pipeline. All code and results were reviewed and verified by the author.